Write each equation in slope-intercept form:
1. 10 - x - 2y = 0
2. y + 8 = -3/4 (x+1)
Determine the equation of each line (write in slope-intercept form)
3. The line through (8, 0) that is perpendicular to the line x + y = 3
4. The line through (0, 6) that is parallel to the line 3x + 5y = 15



Answer by fanism(39)

#1
slope-intercept form = y = mx+c
10 - x - 2y = 0
-2y = -10 +x
y = -(1/2)x + 5
#2
y + 8 = -3/4 (x+1)
y = -(3/4)x - (29/4)